A runner wants to run 10,000m in 30 minutes. After 27 minutes he has run 8,900m. For how long must he accelerates at 0.02 m/s2 and then maintain that final speed


1
Expert's answer
2021-02-22T10:23:12-0500

Let's first find the initial velocity of the runner (before he begins to accelerate):


"v_i=\dfrac{x_i}{t}=\dfrac{8900\ m}{27\ min\cdot\dfrac{60\ s}{1\ min}}=5.5\ \dfrac{m}{s}."

Then, the runner begins to accelerate at a constant acceleration "0.2\ \dfrac{m}{s^2}." We can write the velocity function of the runner as follows:


"v(t)=5.5+0.2t"

We can find the distance at which the runner will accelerate by integrating "v(t)" from 0 to "t":


"d_1(t)=\displaystyle\intop_{0}^tv(t)dt,""d_1(t)=\displaystyle\intop_{0}^t(5.5+0.2t)dt=5.5t+0.1t^2."

Then, we can find the distance that the runner will run in "(180-t)" seconds at constant speed (after he accelerates "t" seconds):


"x_2(t)=(5.5+0.2t)(180-t)."

Finally, adding these two distances, we get:


"x_f-x_i=d_1(t)+d_2(t),""5.5t+0.1t^2+(5.5+0.2t)(180-t)=10000-8900=1100,""0.1t^2-36t+110=0."

This quadratic equation has two roots: "t_1=357\ s" and "t_2=3.1\ s." Since the runner has only 180 seconds to reach the destination, the correct answer is "t=3.1\ s."

Answer:

"t=3.1\ s".